T

he time

that

a laptop battery lasts in everyday use before recharge is needed

is normally

distributed, with a mean of 270 minutes

and

a standard deviation o

f 55 minutes.

a) What is the probability that the battery lasts

more than four hours?

b) What value of

battery

life in minutes is exceeded with 95 % probability?

c) What is the probability

that

the battery lasts exactly

300 minutes?



1
Expert's answer
2022-05-19T11:14:56-0400

Let "X=" the time of a laptop battery: "X\sim N(\mu, \sigma^2 )."

Given "\mu=270min, \sigma=55min."

a)


"P(X>240)=1-P(Z\le \dfrac{240-270}{55})"

"\approx1-P(Z\le-0.54545)\approx0.70728"

b)


"P(X>x)=1-P(Z\le \dfrac{x-270}{55})=0.95"

"\dfrac{x-270}{55}=-1.6449"

"x=270-55(1.6449)"

"x=179.53\ min"

c)


"P(X=300)=0"